OpenStudy (anonymous):

Suppose that r varies directly with s and inversely with t, and r = 2 when s = 3 and t = 12. What is the value of r when s = 5 and t = 4? 2/5 32/5 10 20

OpenStudy (anonymous):

first thing is to solve for K (constant) in the first equation.... r = (ks)/t....so based on given k = 8 now solve r = (ks)/t when s = 5 t =4 and k = 8